A bullet of mass 20kg with a velocity of 15m/s hits a wall and penetrated through a distance of 70cm. what is the bullets retarding force and retardation
1
Expert's answer
2019-06-06T10:00:53-0400

According to Newton's second law, the force "F" acting on something with mass "m" to accelerate/decelerate it with acceleration/retardation "a" correspondingly is


"F=ma."

What is the retardation in our problem?

If the bullet passed 70 cm and decreased its velocity from initial to zero, the retardation is


"a=\frac{v^2_\text{final}-v^2_\text{initial}}{2d}=\frac{0^2-15^2}{2\cdot0.7}=-160.7\text{ m/s}^2."

The negative sign means that the body decelerates. Since we are already talking about retardation, we can ignore it.

The retarding force is


"F=20\cdot160.7=3214.3\text{ N}."